OpenStudy (brittany16):

Solve for the variable. 45 = 30 + m

OpenStudy (lgbasallote):

m = 15

jimthompson5910 (jim_thompson5910):

45 = 30 + m 45 - 30 = 30 + m -30 15 = m m = 15